OpenStudy (linn99123):

need help fast A square-shaped billboard has an area of 295 square feet. Christy wants to put lights around the entire perimeter of the billboard. Approximately how much frame does she need? A. 12 feet B. 17 feet C. 69 feet D. 74 feet

OpenStudy (broskishelleh):

D. would be the closest it is 296 ft perimeter

OpenStudy (austinl):

\(Area = length\times~width\) But, for a square, \(Area = s^2\) Where s is side length. So, we have an area of \(295ft^2\) So putting that in, \(295ft^2=s^2\) Do you think you could solve from there?

OpenStudy (austinl):

For the record, it isn't D... *facepalm*

OpenStudy (broskishelleh):

Ya don't say

OpenStudy (linn99123):

i got a 60 on my test

OpenStudy (anonymous):

I habe the same problem but I don't want to get a bad grade on mine what is the answer

OpenStudy (♥ashley♥):

It's actually C.